A 422 uF capacitor charged to 31.0 Volts discharges completely over 1.51 seconds. What is the time-averaged current, in mA?
Answer #1

C = 422 × 10^-6 F

V = 31V , t = 1.51s

Since Q = CV = 422×10^-6 × 31 = 13082 × 10^-6 C

Now current I = 13.082/1.51 mA

=) I = 8.66mA